Suppose that in a perfectly competitive market, demand is given by Q=59.0-P and supply is given by Q=P-28.0. The government imposes a per-unit excise tax of $1 on the good. What is consumer surplus after that tax is imposed? No units, no rounding.
Answer #1

Answer

after-tax the supply curve is:

Q=P-28

P=Q+28

adding tax

P=Q+28+1=Q+29

P=Q+29

Q=P-29

now the equilibrium at new supply =Qd

P-29+59-P

2P=88

P=44=price consumer pay

Q=59-44=15

Consumer surplus is the area above price and below the demand curve

CS=0.5*(Y-axis intercept of the demand curve -P)*Q

=0.5*(59-44)*15

=112.5

the consumer surplus is 112.5